Han Products manufactures 28,000 units of part S-6 each year for use on its production line. At this level of activity, the cost per unit for part S-6 is:

    

  Direct materials $ 4.20
  Direct labor 7.00
  Variable manufacturing overhead 3.10
  Fixed manufacturing overhead 15.00
  Total cost per part $ 29.30

   

An outside supplier has offered to sell 28,000 units of part S-6 each year to Han Products for $48.50 per part. If Han Products accepts this offer, the facilities now being used to manufacture part S-6 could be rented to another company at an annual rental of $820,600. However, Han Products has determined that two-thirds of the fixed manufacturing overhead being applied to part S-6 would continue even if part S-6 were purchased from the outside supplier.

    

Required:
a.

Calculate the per unit and total relevant cost for buying and making the product? (Round your Per Unit answers to 2 decimal places.)

     

b. How much will profits increase or decrease if the outside supplier’s offer is accepted?
Answer #1

Solution

Han Products

  1. Calculation of per unit and total relevant cost for buying and making the product:

Han Products

Per Unit Differential Costs

28,000 units

Make

Buy

Make

Buy

Cost of purchasing

$48.50

$1,358,000

Cost of making:

Direct materials

$4.20

$117,600

Direct labor

$7.00

$196,000

Variable manufacturing overhead

$3.10

$86,800

Fixed overhead

$5.00

$140,000

Total cost

$19.30

$540,400

Note: avoidable portion of fixed cost is relevant for the decision making.

Avoidable fixed cost = 1/3rd of fixed cost

One third of fixed cost = $15 x 1/3 = $5 per unit

  1. Determination of increase or decrease in profits if outside supplier’s offer is accepted:

Cost of buying S-6 part = $1,358,000

Less: rental revenue from use of released facilities $820,600

Net cost of buying the part S-6 = 1,358,000 – 820,600 = $537,400

Net cost of making the part S-6 = $540,400

Cost of making is higher than cost of buying by $3,000        (540,400 – 537,400)

Hence, profits increase by $3,000 if the company accepts the outside supplier’s offer.